# Semantic Latent Motion for Portrait Video Generation

### v1

@@ 3.2 Motion Generator
In this module, our objective is to train a motion generator that can generate a corresponding motion sequence based on a previous motion sequence, a randomly selected reference motion, and an audio sequence. Our generator is a standard Diffusion Transformer (DiT). We first apply a 0–20% mask Eq. 4 to the previous image feature before extracting the reference motions $\mathbf{M_{p}^{m}}$ and extract target motions $\mathbf{M_{t}}$ without mask Eq. 5, we add noise to the target motions Eq. 6, these are concatenated along the time dimension and processed through self-attention layers. Following this, we use two cross-attention layers: one with the audio feature $\mathbf{A}$ and the other with the reference motion $\mathbf{M_{r}}$, note that the audio feature is embeded by Whisper [27]. In Fig. 9, we demonstrate that applying a mask to the previous image enhances the model’s generalization ability.
$\small\mathbf{M_{p}^{m}}=\operatorname{Enc}(\mathbf{Z^{m}_{p}}\oplus\mathbf{T})$ (4)
$\small\mathbf{M_{t}}=\operatorname{Enc}(\mathbf{Z_{t}}\oplus\mathbf{T})$ (5)
$\small\mathbf{M_{t}^{n}}=(1-t)\mathbf{M_{t}}+t\epsilon$ (6)
$\small\hat{\mathbf{V}}=\operatorname{G}\left(\mathbf{M_{p}^{m}},\mathbf{M_{t}^{n}},\mathbf{M_{r}},\mathbf{A}\right)$ (7)
@@ 3.3 Training Objective
In the first stage Eq. 8, the goal of the motion autoencoder is to reconstruct the target images based on the reference image and corresponding motion. In the second stage Eq. 9, the model is optimized to generate motion sequence based on audio condition. In both stage, the model predicts the noise velocity, and we optimize the mean squared distance between the predicted velocity $v_{\theta}$ and the true velocity $v$.
$\mathcal{L}_{1}=\mathbb{E}_{t,z_{r},z_{t},\epsilon}\left[\left\|(z_{t}-\epsilon)-v_{\theta}\left(z_{r},z_{t},t\right)\right\|^{2}\right]$ (8)
$\mathcal{L}_{2}=\mathbb{E}_{t,m_{r},m_{t},m_{p},\epsilon}\left[\left\|(m_{t}-\epsilon)-v_{\theta}\left(m_{r},m_{t},m_{p},t\right)\right\|^{2}\right]$ (9)
@@ 4.1 Experimental Settings
Figure 6: Qualitative comparison with existing approaches. The results demonstrate that our model effectively generates complex, realistic, and natural motions across various audio conditions.
Table 1: Model Capacity of various methods.
Method Unsupervised Long-term Modeling Real-time Generation Generic Generation Pipeline
Aniportrait [41]
$\times$
$\times$
$\times$
$\times$
Hallo2 [6]
$\times$
$\times$
$\times$
$\times$
SadTalker [48]
$\times$
$\times$
Anitalker [21]
$\times$
Echomimic [4]
$\times$
$\times$
$\times$
$\times$
SeMo (Ours)
Table 2: Motion-based video reconstruction comparison of various methods on the HDTF, DH-FaceVid-1K, and MEAD dataset.
@@
FOMM[30] 24.690 188.797 0.064 25.572 0.816 73.484 30.052 236.578 0.061 25.542 0.844 60.701 41.541 166.950 0.040 31.131 0.917 60.267
MRAA[31] 23.004 155.53 0.046 29.411 0.891 73.653 22.258 216.034 0.044 30.295 0.913 60.481 35.378 167.186 0.030 34.245 0.950 50.558
TPSMM[51] 26.556 138.88 0.041 29.51 0.901 58.478 24.697 215.362 0.042 29.625 0.917 54.784 34.358 136.651 0.028 34.805 0.956 69.224
LIA[39] 26.366 186.307 0.067 26.624 0.840 66.068 33.409 238.101 0.067 26.876 0.861 54.488 47.152 170.724 0.048 31.124 0.921 45.276
Liveportrait[12] 22.469 137.935 0.068 24.810 0.802 59.955 26.502 170.360 0.059 25.681 0.839 52.026 56.700 88.304 0.038 31.661 0.923 44.817
SeMo (Ours) 21.431 136.738 0.043 29.12 0.912 57.849 22.045 112.166 0.036 32.132 0.914 51.437 23.546 70.975 0.019 36.233 0.949 44.517
@@
Implementation Details. All two-stage experiments were trained on 8 NVIDIA A800 GPUs. In Stage 1, we first trained an image-to-image version of the Motion AutoEncoder. Subsequently, we utilized the pre-trained weights to further train it as an image-to-video version. In the main experiments, we maintained the motion representation as 1 token with dimension 512. In Stage 2, we trained the motion generator, generating 16 consecutive motion tokens in a single forward pass. A consistent learning rate of 1e-4 was applied across all experiments.
@@
Baseline. In the motion-based video reconstruction phase, we compared MRAA [31], FOMM [30], Lia [39], Liveportrait [12], TPSMM [51]. In the portrait video generation phase, we compared Aniportrait [41], Anitalker [21], Echomimic [4], Hallo2 [6] , SadTalker [48]
4.2 Comparisons.
Motion-based Video Reconstruction. Motion-based video reconstruction refers to an algorithm that generates target frames based on the reference image and motion. In the experiment, Tab. 2 we directly applied their pretrained weights in the face reconstruction task. We observed that our method outperformed theirs in terms of metrics. Additionally, our approach not only modeled the motion of the face but also modeled the motion in the surrounding areas of the face, making the entire video look more realistic.
Audio-driven Video Generation. Our generation framework achieves remarkably natural motion effects. As shown in Fig. 5 , when processing utterances like ”Yes!”, existing methods only produce limited lip movements, while our approach naturally integrates lip pursing with spontaneous eye blinks to deliver realistic contemplative expressions. Fig. 3 highlights our framework’s robustness across diverse facial poses and camera distances, generating substantial yet plausible motion variations. Notably, our method effectively captures authentic lighting interactions (e.g., dynamic shadow transitions during head movements), significantly enhancing visual realism. Quantitative evaluations in Tab. 4 Tab. 3 confirm our superiority over SOTA methods, with particularly significant improvements in FVD.
Tab. 1 demonstrates the strong capabilities of the model. The framework’s scalability stems from our compact latent motion representation, enabling long-term temporal modeling for holistic event understanding. Through RectifiedFlow’s reflow distillation, we achieve real-time synthesis with minimal sampling steps. As a prior-free and unsupervised architecture, our framework demonstrates strong cross-domain adaptability, enabling efficient application to novel scenarios without architectural modifications.
Table 3: Audio-driven video generation comparison of various methods on the DH-FaceVid-1K dataset.
@@
Table 4: Audio-driven video generation comparison of various methods on the HDTF dataset.
@@
User Study. We randomly selected 10 images from HDTF and DH-FaceVid-1K, and generated 20 videos using 2 random audio clips. We invited 20 participants to choose the best model from 4 different perspectives. We significantly outperform competitors in realism and match the state-of-the-art (SOTA) in lip sync, smoothness, and overall quality.
@@ 4.3 Ablation Study
Dimension and Token Number. In Tab. 5 Tab. 6, we compared the impact of latent motion token dimensions and token numbers on model performance. For the token dimension, we fixed the number at 1 and varied its size, finding that even with a dimension as small as 32, the model could still generate reasonable images. However, smaller dimensions resulted in reduced detailed information, requiring the diffusion model to compensate, which introduced some randomness in the output, though the layout remained consistent. For the token number, we fixed the motion representation space at 512 floats and found that increasing the number of tokens improved performance. This is due to the increased parameters used by the model to analyze motion, enabling more accurate mapping from semantic tokens to low-level pixels.
Table 5: Ablation studies of the latent Motion’s channel dimension on the HDTF dataset.
Token Dimension HDTF
@@
Table 6: Ablation studies of the latent Motion’s token length on the HDTF dataset.
Token Number HDTF

### v2

@@ 3.2 Motion Generator
Let $p$ be the number of previous frames input. In this module, our objective is to train a motion generator that can generate a corresponding motion sequence based on a previous motion $\{m_{\text{pre}}^{f}\}_{f=1}^{p}$, noised target motion $\{\bar{m}_{\text{tar}}^{f}\}_{f=1}^{k}$ and an audio sequence $\{a^{f}\}_{f=1}^{k}$. Note that $\{m_{\text{pre}}^{f}\}_{f=1}^{p}$ and $\{m_{\text{tar}}^{f}\}_{f=1}^{k}$ is extracted by motion encoder from corresponding image features without mask and the audio sequence $\{a^{f}\}_{f=1}^{k}$ is encoded by a pretrained model Whisper [25]. Specifically, we concatenate the noised target motions $\{{\bar{m}}_{\text{tar}}^{f}\}_{f=1}^{k}$ with previous motion $\{m_{\text{pre}}^{f}\}_{f=1}^{p}$ for self-attention modeling, while cross-attention layers establish motion-audio alignment. As shown in fig. 7, the masking strategy improve the distribution of motion latents, which accelerates the convergence as well as the performance of the second-stage training. The motion generation process can be formulated as follows: $\{m_{\text{tar}}^{f}\}_{f=1}^{k}=\operatorname{Gen}_{\phi}(\{\bar{m}_{\text{tar}}^{f}\}_{f=1}^{k},\{m_{\text{pre}}^{f}\}_{f=1}^{p},\{a^{f}\}_{f=1}^{k})$, where $\phi$ is the parameter of the motion generator.
@@ 3.3 Training Objective
In the first stage, the goal of the motion autoencoder is to reconstruct the target video based on the reference image and corresponding motion. In the second stage, the model is optimized to generate motion sequence based on audio condition. Both stage utilize rectified flow as the generative framework. Let $v_{\psi}$ and $u_{\phi}$ denote the velocity respectively, we optimize the distance between the predicted velocity $v_{\theta}$ and the true velocity $v$.
$\min_{\theta,\psi}\mathcal{L}_{\text{first stage}}=\min_{\theta,\psi}\mathbb{E}_{t,x_{\text{ref}},x_{\text{tar}},\epsilon}\left[\left\|(x_{\text{tar}}-\epsilon)-v_{\psi}\left(x_{\text{ref}},\bar{x}_{\text{tar}},\operatorname{Enc}_{\theta}(x_{\text{ref}}),\operatorname{Enc}_{\theta}(x_{\text{tar}}),t\right)\right\|^{2}\right]$ (1)
$\min_{\phi}\mathcal{L}_{\text{second stage}}=\min_{\phi}\mathbb{E}_{{t},a,m_{\text{tar}},m_{\text{pre}},\epsilon}\left[\left\|(m_{\text{tar}}-\epsilon)-u_{\phi}\left(a,\bar{m}_{\text{tar}},m_{\text{pre}},t\right)\right\|^{2}\right]$ (2)
@@ 4 Experiment
Our experiments are mainly divided into three parts. First, we verify the effectiveness of the representations on the video reconstruction task section 4.2.1. Then, we evaluate the applicability of the representations to downstream tasks by testing on audio-driven portrait video generation section 4.2.2. Finally, we conduct extensive ablation studies to demonstrate the necessity of our experimental strategies section 4.3.
@@ 4.1 Experimental Settings
Figure 4: Results of video reconstruction. We show video reconstruction results on datasets DH-FaceVid-1K and HDTF.
Figure 5: Result of Audio-driven Video Generation. Our method adapts to various challenging scenarios. Zoom in for a better view. (a) Close view: the subject is very close to the camera; (b) Distant view: the subject is far from the camera, nearly half-body is shown; (c) Edge view: the face appears at the edge of the frame; (d) Partial face: only part of the face is visible; (e) Head turn: the subject turns their head.
Table 1: Motion-based video reconstruction comparison of various methods on the HDTF, DH-FaceVid-1K, and MEAD dataset. Red indicates the best result, blue the second best, and green the third best.
@@
FOMM[29] 24.690 188.797 0.064 25.572 0.816 30.052 236.578 0.061 25.542 0.844 41.541 166.950 0.040 31.131 0.917
MRAA[30] 23.004 155.53 0.046 29.411 0.891 22.258 216.034 0.044 30.295 0.913 35.378 167.186 0.030 34.245 0.950
TPSMM[55] 26.556 138.88 0.041 29.51 0.901 24.697 215.362 0.042 29.625 0.917 34.358 136.651 0.028 34.805 0.956
LIA[39] 26.366 186.307 0.067 26.624 0.840 33.409 238.101 0.067 26.876 0.861 47.152 170.724 0.048 31.124 0.921
Liveportrait[11] 22.469 137.935 0.068 24.810 0.802 26.502 170.360 0.059 25.681 0.839 56.700 88.304 0.038 31.661 0.923
SeMo (Ours) 21.431 136.738 0.043 29.12 0.912 22.045 112.166 0.036 32.132 0.914 23.546 70.975 0.019 36.233 0.949
@@
Implementation Details. All two-stage experiments were trained on 8 NVIDIA A800 GPUs. We set the dimension of the semantic latent motion to 1 token with 512 channels throughout all experiments to balance reconstruction accuracy and generation quality. In the first stage, the motion autoencoder generates a sequence of $k=16$ frames in a single forward pass. In the second stage, the motion generator also produces $k=16$ motion tokens per forward pass aided with $p=8$ previous motion tokens. All models are trained under similar settings: a base learning rate of $10^{-4}$ per 64 batch size and the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\text{decay}=0.05$.
@@
Baseline. In the motion-based video reconstruction phase, we compared MRAA [30], FOMM [29], Lia [39], Liveportrait [11], TPSMM [55]. In the portrait video generation phase, we compared Aniportrait [42], Anitalker [19], Echomimic [4], Hallo2 [6], SadTalker [52].
4.2 Comparison
4.2.1 Motion-based Video Reconstruction.
Motion-based video reconstruction refers to an algorithm that generates target frames based on the reference image and motion. In the experiment, table 1 we directly applied their pretrained weights in the face reconstruction task. We observed that our method outperformed theirs in terms of metrics. Additionally, Since the training task involves reconstructing entire videos through self-supervised learning rather than focusing solely on the face, in fig. 4,we observe that the model is able to capture more detailed and complex motions, such as those of earrings and hairstyles. Furthermore, thanks to the powerful generative capability of the motion decoder, the model can automatically complete missing details caused by head movements.
Methods based on explicit representations [29, 30, 55] perform poorly, especially in cases involving large or rapid motions. They often fail to capture fine-grained, high-frequency changes. Moreover, these explicit representations are typically designed for the face only, neglecting global motion information, which further limits their performance and robustness. As a result, such methods tend to generalize poorly and cannot be effectively transferred to other scenarios. We further provide extensive insights on the mask section D.1 and scene transferability section D.4.
Figure 6: (a) Comparison of efficiency and quality among audio-driven video generation methods. Bits per Pixel is calculated as $\frac{S\cdot D}{256^{2}}$, where $S$ is the representation size and D is the bit width of each element, as proposed by  [28]. (b) User study on audio-driven video generation. The detail procedure of user study can be found in appendix C.
Table 2: Video generation comparison on DH-FaceVid-1K and HDTF.
@@ 4.2 Comparison
4.2.2 Audio-driven Video Generation.
The quantitative metrics for audio-driven video generation are presented in table 2, while the qualitative (visualization) results are shown in fig. 5. The Sync-C and Sync-D metric was not evaluated on the DH-FaceVid-1K dataset because the models used for evaluation were not trained on Chinese datasets. Our method not only achieves highly realistic visuals, but also excels in robustness across diverse scenarios. As shown in fig. 5, it effectively handles challenging cases where others often fail. Competing methods lack eyeball detail in close-ups, produce unnatural head turns, and frequently fail or generate unrealistic content in partial-face situations. These limitations point to a common issue: prior-based methods [42, 6, 7, 4] require strict assumptions about scene conditions, making such methods heavily scenario-dependent. Typically, they rely on complex data pre-processing pipelines to tightly crop and center the face within the frame. In contrast, our self-supervised approach removes the dependency on prior knowledge, enabling robust performance in unconstrained settings. As shown in fig. 6(b), our user study results demonstrate that our method achieves an impressive 81 % win rate for perceived realism. This indicates that our approach captures more nuanced facial motions, while others often produce noticeable rigid-body translations—a subtle difference easily perceived by human evaluators but not always captured by standard quantitative metrics.In fig. 6(a), we compare the encoding efficiency of each method, which is calculated as the size of the latent representation per frame in the latent space divided by the number of pixels per frame. More visual results can be found in section D.3 and video materials.
@@ 4.3 Ablation Study
Table 3: Ablation studies on the number of motion tokens and the token channel dimension, where T denotes the number of tokens and D denotes the channel dimension.
Token HDTF